I would appreciate help on this question. I answered it but it is still wrong if i can get any help I would really appreciate it. The question is: "A rocket is fired upward from some initial distance above the ground. its height in feet, h, above the ground, t, seconds after it is fired, is given by h= -16t^2+176t+1632.
1.What is the rockets maximum height?
2.How long does it take for the rocket to reach its maximum height?
3.After it is fired, the rocket reaches the ground at t=___?"

Answer by Alan3354(69443)

The question is: "A rocket is fired upward from some initial distance above the ground. its height in feet, h, above the ground, t, seconds after it is fired, is given by h= -16t^2+176t+1632.
---------------
1.What is the rockets maximum height?
Max height is the vertex of the parabola, at t = -b/2a
t = -176/-32 = 5.5 seconds
h(5) = -16*30.25 + 176*5.5 + 1632
h(5) = 2116 feet
----------------------------
2.How long does it take for the rocket to reach its maximum height?
5.5 seconds
-----------------------------
3.After it is fired, the rocket reaches the ground at t=___?"
h= -16t^2+176t+1632 = 0
-t^2 + 11t + 102 = 0
t = -6 (Ignore)
t = 17 seconds